Checking to see if BS01 still have the UV ghosting seen in the in-water and predeployment data
BS03cfg008 = BS01 out of housing with no shtr block aor falcon controler
All the raw data files and their images , Table of KEYWORDS variable |
Page Number |
Link |
Description |
Date |
1.01 |
|
DAY01 - Looking for ghosts with lamp and filters on Track 1 - 450 nm LP filter |
30 Jan 2018 |
1.02 |
|
DAY02 - Looking for ghosts with lamp and filters on Track 3, 5, 7, 9, 11 and 13 - 450 nm LP filter |
30 Jan 2018 |
1.03 |
|
DAY03 - Looking for ghosts with lamp and filters on Track 1, 3, 5, 7, 9, 11 and 13 - Short Pass Fi;ter |
30 Jan 2018 |
1.04 |
|
DAY04 - solar data onto track 5 - ghosting with Es and fraun lines |
7 Feb 2018 |
1.05 |
|
DAY05 - solar data onto track 5 - fiber bends + vertical spec |
7 Feb 2018 |
|
Day01 - Looking for ghosts with lamp and filters on Track 1 |
2.01 |
UV ghosting |
Checking track 1 for UV ghosts |
30 Jan 2018 |
Day02 - Looking for ghosts with lamp and filters on Track 3, 5, 7, 9, 11 and 13 |
3.01 |
UV ghosting |
Checking track 3, 5, 7, 9, 11 and 13 for UV ghosts |
30 Jan 2018 |
Day03 - Looking for ghosts with lamp and filters on Track 1, 3, 5, 7, 9, 11 and 13 - Short Pass Fi;ter |
4.01 |
UV ghosting |
Checking track 1, 3, 5, 7, 9, 11 and 13 for UV ghosts |
30 Jan 2018 |
Day04 - Still looking at UV ghosts but for Track 5 only and using the Solar lamp with fiber changes |
5.01 |
UV ghosting |
Graphs of the ghosting from track 5 onto Track 10 with Solar input |
7 Feb 2018 |
5.02 |
Fiber "modes" |
First fiber "mode" test w/ and w.o fiber and with bends. |
8 Feb 2018 |
5.03 |
Comp Uv ghosts |
Comparing the UV ghosts shapes from day 1 - 5 |
12 Feb 2018 |
Day05 - solar data onto track 5 - fiber bends + vertical spec |
6.01 |
Fiber "mode" & vert |
Mike doing more fiber "modes" testing and one data set with BS veritcal |
9 Feb 2018 |
Compilation of day data |
20.01 |
UV nets->BS link |
RS pow lix data are negative- We found some nevative values in the low pixel on H19-01 and are checking we see them here |
31 Jan 2019 |
DAY 1 EMAIL from Mike |
Hi Steph,
On Friday, after the telecon I went looking for our long-pass and short-pass filters,
to use to look for the ghost in post-M262 BS01 - as per Mike Kehoe's email on 25Jan2018.
I found the filter-holders for Semrock LP & SP filters (frm BCJ?) we had used previously,
but I could not locate the filters. I did find and OLD set of filters Dennis purchased in 1996,
which had a possibly-appropriate 1 in dia long-pass filter = Oriel 52806 = 450 nm LP.
This part number did not show up in I google searches, but I did find an ca. 1990 Oriel catalog,
and I found some papers with that came with the original DKC purchase.
So, on Friday, I setup the BS01cfg008 (no shutter block, no Falcon controller) and
illuminated only fiber bundle FO#1 with the Oriel incandescent lamp via the Spectralon sphere
then took scans without and with the 450nm LP filter. I took bac scans to match the sig scans.
Then I added the 450nm LP filter between the lamp and the sphere and took bac/sig scans.
At the end I turned off the lamp and took some drk scans.
I looked at the data this morning (Monday) and I can BARELY see a ghost sig @ Trk#14!
You must have VERY good eyes!
I don't see any difference in the without-filter vs with-filter net signals past trk pix 200,
but maybe you will?
I think I'll step this no-filter vs with-filter scanning through the other tracks today.
Late Friday I found the Semrock filters (!), so that might also get done today (?).
Day01 =
5x Oriel filter doc scans
1x log sheet scans
5x pairs of sig & bac with no filter
5x pairs of sig & bac with filter
5x drk scans
/ftp1/Mike/L267/bs1cal/doc
/ftp1/Mike/L267/bs1cal
'Njoy! MF
|
DAY 2 EMAIL from Mike |
Hi Steph,
I hope you had not yet started to work on BS01 day01...
because I forgot to make the subdir .../day01/ (now I am having trouble moving the day01 .fits into it...),
and I renamed the day01 log sheet (dough).
The day02 data are T3,5,7,9,11,13 each individually lit via no-filter, then with-filter=Oriel-52086
each with 5x sig,bac pairs (i.e. no-filt & with-filt x 6 Trks = 60x sig + 60x bac),
then 5x drk scans @ end = lamp off, T13 @ sphere.
1x new logsheet @
/ftp1/Mike/L267/bs1cal/doc/
4x photos of day02 setup @
/ftp1/Mike/L267/bs1cal/photos/
125x *.fits files @
/ftp1/Mike/L267/bs1cal/day02/
Luck! MF
P.S. I was just tinking: the drk scans might not tell us everyting I was hoping for,
because the lamp was turned off, so the extra ambient lighting on the
covered fibers not-at-the-sphere-exit-port-jig is less than during the sig scans w/ the lamp on...
I was concerned that the VERY LITTLE not-Trk#1 sig I was seeing @ day01
might have been leaking into the covered fibers not-at-the-sphere...
With the lamp off, the only possible ambient light leak would be via the
acquisition laptop and a desk light...
Either way though, it was pretty dark at the BS...
|
DAY 3 EMAIL from Mike |
Hi Steph,
There are some more BS01 data for you at the NAS please.
Tonight I ran bare-lamp and a short-pass filter via T1,3,5,9,11,13 (i.e. skip T7),
where the SP filter was Semrock FF01-720/SP-25, from BCJ -
there are some Semrock docs in the dir ...bs1cal/doc/.
I did the 5x drk scans at the end a little different tonight =
left the lamp on and capped-&-black-cloth-covered all FO tips...
which still looked dark to me!
There are 125x .fits files @
/ftp1/Mike/L267/bs1cal/day03/
1x new log sheet & ?x new filter docs @
/ftp1/Mike/L267/bs1cal/doc/
Thank you, MF |
DAY 4 EMAIL from Mike |
Hi Steph,
Today I got a few sun scans with the BS01cfg008 Trk#5,
for more ghost-looking with a UV-rich source, plus wavelength kine.
I mirrored sun light into the Spectralon sphere, and
got 5x sig,bac pairs with the bare bundle FO #5 at the sphere exit port.
Then I added a 50 m long 600 um fiber between bundle FO#5 & the sphere
(this was an backup/unused(?) FO from the FISH) and got another 5x sig,bac pairs.
Then I had reposition the long FO to even out a loop of fiber, and got another 5x sig,bac pairs.
At the end I covered the entrance port to the sphere,
leaving Trk#5 FOs connected to the exit port, and collected 5x dark (via sig) scans
There are:
1x new log sheet @ /ftp1/Mike/L267/bs1cal/doc/
20x new photographs @ /ftp1/Mike/L267/bs1cal/photos/
35x fits files @ /ftp1/Mike/L267/bs1cal/day04/
I had to made a rack to hold the BS vertically, with the FO bundle pointing down,
as is it's orientation in a deployed buoy, so that took up the remainder of I day.
I took some photos of this new setup.
Tomorrow, I plan to try bending the long FO near the BS and near the sphere,
with the BS horizontal, then try tipping the BS vertical and repeating,
with sunlight as the source.
Aloha, MF
|
DAY 5 EMAIL from Mike |
Hi Steph,
I am most curious to look at the BS01 day04 data you just processed, but first...
Today/Thursday I ran the BS01 again with the long fiber via the sphere + sun = day05.
Set#1= 5x sig/bac pairs with the long fiber between BS trk#5 & Spectralon sphere.
Set#2= 5x pairs after disconnecting and re-connecting the long fiber
(because I thought it would be safer to tip the BS with the FO disconnected).
Set#3= 5x pairs with a bend added to the long FO near the BS/output end.
Set#4= 5x pairs with the 1st bend removed, then another FO bend added near the sphere/input end.
Set#5= 5x pairs with both bends removed = back to same cfg as set#2.
Set#6= 5x pairs with the BS vertical, after dis-connecting the long FO to move the BS,
then reconnecting the long FO ... only at the BS end = sphere end left alone.
I took photos and tried to name them with the set#'s above.
1x logsheet @ /ftp1/Mike/L267/bs1cal/doc/
30x s,b pairs @ /ftp1/Mike/L267/bs1cal/day05/
21x photos @ /ftp1/Mike/L267/bs1cal/photos/
Thank you, MF
|